esther12345:

If a sphere’s volume is doubled, what is the corresponding change in its radius? A. The radius is increased to 2004-06-01-05-00_files/i0280000.jpg times the original size. B. The radius is increased to 4 times the original size. C. The radius is increased to 2 times the original size. D. The radius is increased to 8 times the original size.

jhonyy9:

do you know formula for a sphere's volume ?

jhonyy9:

asker is offline

jhonyy9:

@ThisGirlPretty

sillybilly123:

\(V =\dfrac{4}{3} \pi r^3\) \(V' =\dfrac{4}{3} \pi r'^3 \) \(\dfrac{V'}{V} = \dfrac{\dfrac{4}{3} \pi r'^3}{\dfrac{4}{3} \pi r^3} = \left(\dfrac{ r'}{ r}\right)^3 = 2\) \(\dfrac{ r'}{ r} = 2^{\frac{1}{3}}\)

sillybilly123:

@jhonyy9 is a troll in case you haven't worked that out ;)

jhonyy9:

@sillybilly123 r' = 2r yes ?

sillybilly123:

good luck QC !!

jhonyy9:

@ThisGirlPretty do you understand @sillybilly123"s work ?

Vocaloid:

the missing value provided in the question is \[\sqrt[3]{2}\]